Patricia wants to invest a sum of money today that will yield $17,000 at the end of 6 years. Assuming she can earn an interest rate of 6% compounded annually, how much must she invest today?
Answer #1

FV = PV (1+r)^n

FV = Future Value, ie, (investment value + yeild)

PV = Investment value today

r = Interest rate

n = No of years

yield = 17000

FV = PV + yield ie PV+17000

FV = PV (1+r)^n

PV+17000 = PV (1+r)^n

(PV(1+.06)^6) - PV = 17000

1.4185 PV - PV = 17000

0.4185 PV = 17000

PV= 17000/0.4185

PV= 40619.41

INVESTMENT AMOUNT TODAY = 40619.41